OpenStudy (anonymous):

Medal: Write an equation in slop intercept form of the line through points (-2,6) and (7,-1)

OpenStudy (utterly_confuzzled):

(-2,6) (7,-1) Y-y1 = m(x-x1) M = slope = -7/9 y - 6 = (-7/9)(x+2)

OpenStudy (utterly_confuzzled):

And solve for y from there

OpenStudy (utterly_confuzzled):

Sorry, I was thinking point-slope, not slope intercept But just take the previous equation and solve for y and that should give you your answer